### v1

# Regularized Least Squares ERM Empirical Risk Minimization

Abstract
We explore the low-rank structure of the weight matrices in neural networks originating from training with Gradient Descent (GD) and Gradient Flow (GF) with $L2$ regularization (also known as weight decay). We show several properties of GD-trained deep neural networks, induced by $L2$ regularization. In particular, for a stationary point of GD we show alignment of the parameters and the gradient, norm preservation across layers, and low-rank bias: properties previously known in the context of GF solutions. Experiments show that the assumptions made in the analysis only mildly affect the observations.
In addition, we investigate a multitask learning phenomenon enabled by $L2$ regularization and low-rank bias. In particular, we show that if two networks are trained, such that the inputs in the training set of one network are approximately orthogonal to the inputs in the training set of the other network, the new network obtained by simply summing the weights of the two networks will perform as well on both training sets as the respective individual networks. We demonstrate this for shallow ReLU neural networks trained by GD, as well as deep linear and deep ReLU networks trained by GF.
1 Introduction
First-order optimization algorithms, such as Stochastic Gradient Descent (SGD) have emerged as a go-to tool for training machine learning models. Their popularity primarily stems from good scalability and empirical performance, while their theoretical properties and performance are reasonably well understood for learning problems with sufficient structure, such as convexity or even weak forms of non-convexity. However, in the recent years with the advent of overparameterized deep neural networks, there was a surge of interest in understanding the behavior of these algorithms on non-convex and often non-differentiable problems. Here, a combination of neural network architecture choice, initialization, and hyperparameter tuning often achieves near-zero training loss while enabling good test-time generalization ability.
Recently, some progress was made in attempt to explain this phenomenon by looking for implicit biases in the algorithm (Vardi, 2023). While implicit biases were known before in the context of simpler learning problems such as overparameterized linear least-squares (solving this problem through pseudo-inverse gives an interpolating solution with the minimal $L2$ norm), observing the same phenomenon in neural network learning is rather recent. Lyu and Li (2020), Ji and Telgarsky (2020) showed that for a certain family of neural networks (positive homogeneous neural networks, such as deep linear or deep ReLU neural networks) trained by an idealized version of Gradient Descent (GD) known as Gradient Flow (GF), asymptotically converges to the predictor with the smallest parameter $L2$ norm. In another influential line of work, known as the Neural Tangent Kernel (NTK) approximation Jacot et al. (2018), Du et al. (2018b), Ji and Telgarsky (2019b), Arora et al. (2019), it was shown that a sufficiently wide shallow non-linear neural network behaves similarly as a linear predictor on Reproducing kernel Hilbert space (RKHS). This enabled reduction to analysis of GD on RKHS, which is known to converge to the minimum $L2$ norm interpolating solution (or, which behaves as a regularized solution when stopped early (Yao et al., 2007)).
While $L2$ norm minimization (or regularization) has a clear interpretation in linear models, its role in deep neural network learning is less intuitive. On that front, several works showed that this form of regularization induces a low-rank structure in weight matrices of neural networks. In fact, in deep linear neural networks weight matrices in such interpolants are shown to be rank-$1$ matrices (Ji and Telgarsky, 2019a). Similarly, Timor et al. (2023) showed that weight matrices in minimum $L2$ norm interpolating deep ReLU networks will have a low stable rank (ratio between Frobenius and spectral norms) as long as the depth is sufficiently large. Phuong and Lampert (2021) show convergence of a hidden weight matrix in a shallow neural network to a rank-one matrix when it is trained by GF on orthogonally-separable data (in the context of classification all inputs with a matching label satisfy $\left\langle x_{i},x_{j}\right\rangle>0$, and otherwise for a non-matching one). Min et al. (2024) show convergence of stable rank to $2$ for shallow ReLU networks trained by GF dynamics. Under mild assumptions on the inputs Frei et al. (2023) established that the hidden weight matrix of a shallow neural network convergences to the low-rank matrix (at most $2$).
The presence of these biases was established for training without any form of explicit regularization, however one might suspect that a similar message can be said when we introduce explicit regularization. Some recent works have indeed investigated whether a low-rank structure arises in stochastic optimization (Galanti et al., 2024), with the emphasis on the role of the batch size. Others have explored $L2$ regularization from a generalization viewpoint, in the NTK approximation setting, when neural networks behave similarly to overparameterized linear models (Wei et al., 2019, Hu et al., 2021).
Finally, while the appearance of low-rank weight matrix structure during training is interesting on its own, here we might ask a related question whether low-rank bias can explain other phenomena that we observe in neural network learning. In particular, in this paper we look at the connection between low-rank bias and a form of a multi-task learning known as model merging. In model merging, weight matrices of two neural networks trained on different tasks are summed to form a new neural network, which often performs well on both original tasks. This behaviour seems surprising at first glance, and the reason why merging (or, related, parameter averaging) in neural networks might be effective is not well understood.
1.1 Our contributions
In this paper we take a closer look at a low-rank bias and its effect on model merging. Our contribution is two-fold yet at the core it is connected to low-rank bias induced by $L2$ regularization.
Model merging enabled by weight decay.
We show that low-rank bias in neural networks enables effective model merging: after neural networks are trained on different datasets with mutually nearly-orthogonal inputs (but not necessarily orthogonal within the task), simply summing their weight matrices results in a predictor with combined weights that performs well simultaneously on all tasks. We explain this phenomenon through the low-rank bias. Training biases different networks toward low-rank weight matrices that span non-overlapping subspaces. Consequently, summing their weight matrices results in a neural network that behaves as the original ones on inputs from respective tasks. In other words, two neural networks can be made to ‘reside’ in one parameter set.
More formally, given input-label pairs $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\in\mathbb{S}^{d-1}\times\{\pm 1\}$ we obtain a predictor $f_{\theta(t)}$, whose parameters $\theta(t)$ are found by GD (or GF) run for $t$ steps, while minimizing a regularized loss $\theta\mapsto\frac{1}{n}\sum_{i}\mathrm{loss}(f_{\theta}(x_{i}),y_{i})+\lambda\|\theta\|^{2}$. In a similar way, we obtain parameters $\theta^{\prime}(t)$ given another data $(x_{1}^{\prime},y_{1}^{\prime}),\ldots,(x_{n}^{\prime},y_{n}^{\prime})$ belonging to the second task, such that inputs between these tasks are approximately orthogonal in a sense that $\max_{i,j}\left|\left\langle x_{i},x_{j}^{\prime}\right\rangle\right|\leq\varepsilon$. Then, given an input $x$ originating from the first task, we show that
$|f_{\theta(t)}(x)-f_{\theta(t)+\theta^{\prime}(t)}(x)|\leq f_{\theta^{\prime}(0)}(x)\,e^{-\lambda t}+C_{2}\,\varepsilon$
for several scenarios. Here, exponentially vanishing term that appears because of $L2$ regularization, is responsible for contribution of initialization, which is typically non-zero in neural network training. The second, $\varepsilon$-dependent term captures the length of projection of the input (or activation vector in case of multilayer neural networks) from one task onto the weight matrix of the network trained on another task. Note that $L2$ regularization is essential here, since without it, the effect of initialization (appearing through a constant term $f_{\theta^{\prime}(0)}(x)$) would not disappear. Interestingly, none of these results require convergence to the local minimum.
In particular, we show the above in case of linear prediction and shallow ReLU neural networks trained by GD, and deep linear neural networks trained by GF. In case of deep ReLU neural networks we look at a simplified case where $\varepsilon=0$ and $t\to\infty$. In that case we conclude that $f_{\theta(t)}(x)=f_{\theta(t)+\theta^{\prime}(t)}(x)$. In this case we only require that GF reaches a stationary point. Finally, in Section 6.2 we provide empirical verification for these results for fully-connected ReLU neural networks, which to some extent corroborate the above.
Low-rank bias and alignment through weight decay
Model merging discussed earlier crucially relies on $L2$ regularization, and its success (good performance on both tasks) can be attributed to biases that arise in neural networks because of regularization. To this end, in Section 4 we take a closer look at such biases. In particular, this time we focus on homogeneous neural networks with Lipschitz-gradient activations (such as, e.g. ReLUp with $p\geq 2$), and consider running GD with $L2$ regularization until convergence to the stationary point. We first establish asymptotic alignment, meaning that parameters converge in the direction of gradient of the (unregularized) loss. Although asymptotic alignment was known before in case of GF (without regularization) (Ji and Telgarsky, 2019a, 2020), here we observe it in case of GD. Building upon this observation we prove several other results that, which to the best of our knowledge, were only known in case of GF (Du et al., 2018a, Ji and Telgarsky, 2019a, 2020, Timor et al., 2023):
  • Norm-preservation: weight matrices in neural networks we consider have the same Frobenius norm throughout layers.
  • Deep linear neural networks asymptotically have all weight matrices of rank one.
  • Deep (possibly non-linear) neural networks have a low-rank weight matrix structure, in the following sense:
The harmonic mean of stable ranks [^1] A stable rank of matrix is defined as a ratio of its Frobenius and spectral norms, while it is small whenever matrix has few dominant eigenvectors. of weight matrices $(W_{1},\ldots,W_{K})$ is controlled by regularization parameter $\lambda$:
$\frac{K}{\left(\frac{\|W_{1}\|_{F}}{\|W_{1}\|_{2}}\right)^{-1}+\dots+\left(\frac{\|W_{K}\|_{F}}{\|W_{K}\|_{2}}\right)^{-1}}<\sqrt{\frac{L(\theta)^{\frac{1}{2}-\frac{1}{K}}}{\lambda}}\leq C\,\lambda^{\frac{1}{K}-1}$
Harmonic mean of stable ranks as a proxy for capturing the rank structure in deep neural networks was proposed by Timor et al. (2023). Since stable rank cannot be smaller than one, the harmonic mean is small whenever majority of weight matrices have small stable ranks. In particular, Timor et al. (2023) looked at the minimum $L2$-norm interpolating neural networks, in a sense that for all training examples $y_{i}\in\{\pm 1\}$, $y_{i}f_{\theta}(x_{i})\geq 1$. Their conclusion was that the harmonic mean $\to\sqrt{2}$ as depth $K\to\infty$. In contrast, we simply require GD to achieve a stationary point, in which case the harmonic mean is controlled by regularization parameter and the training loss at stationarity, even at a finite depth (see Lemma 4 and discussion therein). Indeed, this seems to be supported by some basic empirical evidence (see Sections 6.1 and 2), which suggests that that a low stable rank can be achieved even without interpolation but with weight decay.
Finally, we look at the behavior of a stable rank and norm preservation empirically for very deep neural networks (such as pre-trained large language models) in Section 6.1, and conclude that assumptions (such as homogeneity and stationarity) only mildly affect the observations.
2 Definitions
[…]
In the following we denote Rectified Linear Unit (ReLU) operation by $(x)_{+}=\max(x,0)$ for $x\in\mathbb{R}$. For vectors and matrices application of $(\cdot)_{+}$ is understood elementwise. ReLU has several useful properties. For $a,b\in\mathbb{R}$, we have $|(a)_{+}-(b)_{+}|\leq|a-b|$ and so for vectors $x,y\in\mathbb{R}^{d}$, $\|(x)_{+}-(y)_{+}\|_{2}^{2}=\sum_{i}((x_{i})_{+}-(y_{i})_{+})^{2}\leq\sum_{i}(x_{i}-y_{i})^{2}=\|x-y\|_{2}^{2}$.
Function $f:\mathbb{R}^{d}\to\mathbb{R}$ is called positive homogeneous of degree $K$ when $f(\alpha x)=\alpha^{K}f(x)$ for any $\alpha\geq 0$. An important consequence for such functions is an Euler’s homogeneous function theorem which states that, if the chain rule holds, then $Kg(x)=\left\langle x,\nabla f(x)\right\rangle$. Note that ReLU is positive homogeneous, meaning that $(\alpha x)_{+}=\alpha(x)_{+}$ for $\alpha\geq 0$. In particular, this implies that for a $K$-layered ReLU or linear neural network is positive homogeneous and so $f_{\theta}(\alpha\,x)=\alpha^{K}f_{\theta}(x)$ and so Euler’s theorem holds.
For the logistic loss function $\ell(x)=\ln(1+e^{-x})$ we have $\ell^{\prime}(x)=-e^{-x}/(1+e^{-x})\in(-1,0)$ and moreover $|\ell^{\prime}(x)|=-\ell^{\prime}(x)\leq\ell(x)$ and $x\,\ell^{\prime}(x)\leq\ell(x)$ for all $x$.
Differentiation
[…]
$\partial F(\theta)=\mathrm{conv}\left(\left\{g\in\mathbb{R}^{p}:\exists(\theta_{i})_{i}\to\theta,\quad\nabla F(\theta_{i})\to g\right\}\right).$
Vectors in $\partial F$ are called subgradients, while ¯$\bar{\partial}f(x)$ will stand for a unique minimum-norm subgradient, that is ¯$\bar{\partial}f(x)=\operatorname*{arg\,min}_{g\in\partial f(x)}\|g\|~{}$. Throughout the paper it is understood that $\nabla f=\bar{\partial}f$. We will assume that the chain rule holds, that is $\dot{f}(\theta(t))=\langle g(t),\dot{\theta}(t)\rangle,$ for all $g(t)\in\partial F(\theta(t))$. It is possible to formally establish that the chain rule holds by assuming a technical notion of ‘definability’ (not covered here, see (Ji and Telgarsky, 2019a)), which excludes functions that might result in a badly behaved optimization.
3 Preliminaries
In this paper we consider multilayer neural networks $f_{\theta}:\mathbb{R}^{d}\to\mathbb{R}$ given by a recursive relationship
$f_{\theta}(x)=\left\langle w_{K},h_{K-1}\right\rangle,\quad h_{k}=\sigma(W_{k}h_{k-1}),\quad h_{0}=x\qquad(k\in[K-1])$
where $x$ is the input, $(W_{k}\in\mathbb{R}^{m\times m})_{k}$ is collection of weight matrices, $\theta=(\mathrm{vec}(W_{1}),\ldots,\mathrm{vec}(W_{K}))$ is a parameter vector, and $\sigma:\mathbb{R}^{m}\to\mathbb{R}^{m}$ is activation function. When we look at linear neural networks $\sigma$ is identity, meanwhile in the non-linear case or ReLU networks $\sigma(x)=(x)_{+}$. Activation vectors can alternatively be written as
$h_{k}=D_{k}W_{k}D_{k-1}W_{k-1}\cdots D_{1}W_{1}x$
where $D_{k}$ is a diagonal matrix $D_{k}=\operatorname*{diag}(\sigma^{\prime}(W_{k}h_{k-1}))$ while $(D,h)$ implicitly depend on the input $x$. When we want to state an explicit dependence on the inputs and parameters we will use notation $h_{k}(x;\,\theta)$. Then, differentiation with respect to a weight matrix has a convenient form
$\frac{\mathop{}\!\mathrm{d}f_{\theta}(x)}{\mathop{}\!\mathrm{d}W_{k}}$ $=(W_{K}D_{K-1}\cdots W_{k+1}D_{k})^{\top}(h_{k-1})^{\top}~{}.$
In practice parameters $\theta$ are tuned based on the training data by minimizing some loss function. In this paper we will focus on a binary classification problem, and so given a tuple of inputs and labels $(x_{i},y_{i})_{i=1}^{n}\in(\mathbb{B}^{d}\times\{-1,1\})^{n}$, the training loss (or empirical risk) is defined as
$L(\theta)=\frac{1}{n}\sum_{i=1}^{n}\ell\left(y_{i}\,f_{\theta}(x_{i})\right)$
with respect to a logistic loss function $\ell(x)=\ln(1+e^{-x})$. More generally we will look at the regularized objective of a form
$L_{\lambda}(\theta)=L(\theta)+\frac{\lambda}{2}\|\theta\|_{2}^{2}\qquad(\lambda\geq 0)~{}.$
GD algorithm approximately minimizing $\theta\mapsto L_{\lambda}(\theta)$ is given by Algorithm 1.
Algorithm 1 Gradient Descent
1:  Input: step size $\eta>0$, data $(x_{i},y_{i})_{i}$, initial parameter $\theta_{0}\in\mathbb{R}^{p}$.
2:  for $s=0,1,2,\ldots,t$ do
3:     For all $k\in[K]$:
$W_{k,s+1}=(1-\eta\lambda)W_{k,s}-\frac{\eta}{n}\sum_{i=1}^{n}y_{i}\,\ell^{\prime}(y_{i}\,f_{\theta_{s}}(x_{i}))\,\frac{\mathop{}\!\mathrm{d}f_{\theta}(x_{i})}{\mathop{}\!\mathrm{d}W_{k}}(\theta_{s})$
4:  end for
5:  Output $\theta_{t}=(\mathrm{vec}(W_{1,t}),\ldots,\mathrm{vec}(W_{K,t}))$.
When we consider GF dynamics (we use $(t)$ time indexing instead of $\cdot_{t}$ as in the discrete case), the update rule is replaced by time derivative
$\dot{W}_{k}(t)=-\lambda W_{k}(t)-\frac{1}{n}\sum_{i=1}^{n}y_{i}\,\ell^{\prime}(y_{i}\,f_{\theta(t)}(x_{i}))\,\frac{\mathop{}\!\mathrm{d}f_{\theta}(x)}{\mathop{}\!\mathrm{d}W_{k}}(\theta(t))~{}.$ (1)
4 Stationary points and low-rank bias
In this section, we study training deep neural networks with gradient descent and weight decay. We first consider the case of differentiable objective with Lipschitz gradient. Examples of such are neural networks with smooth activation functions (such as identity or smoothed versions of ReLU, e.g. powers of ReLU).
Lemma 1 (Alignment).
Assume $L_{\lambda}$ is differentiable and $H$-smooth with $H>0$ ($\nabla L_{\lambda}$ is $H$-Lipschitz w.r.t. $L2$ norm) and that step size satisfies $\eta\leq 1/H$. Then, the limiting point $\theta=\theta_{\infty}$ satisfies $\lambda\theta=-\nabla L(\theta)$. [Proof]
This type of alignment was first studied by (Ji and Telgarsky, 2019a, 2020) in the context of training with gradient flow, and under the extra condition that $L(\theta_{0})<\ell(0)$. Here, weight decay enables a simpler argument to establish alignment. Next, we show several implications of this result.
Lemma 2 (Deep linear networks).
Let $f_{\theta}$ be a deep linear network, and $\theta$ be a parameter vector satisfying the alignment condition of Lemma 1. Then all weight matrices $W_{1},\dots,W_{K-1}$ are rank-1. [Proof]
Lemma 3 (Norm preservation).
Let $\theta$ be a parameter vector satisfying the alignment condition of Lemma 1. Further, assume that $f_{\theta}$ is locally Lipschitz and positively homogeneous. Then weight matrices of all layers have the same Frobenius norm:
$\lambda\|W_{k}\|_{F}^{2}=-\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime}(y_{i}f_{\theta}(x_{i}))y_{i}f_{\theta}(x_{i})\;.$
[Proof]
A deep neural network with non-linear smooth activations can satisfy the above conditions. A self-attention layer used in Transformer architecture (Vaswani et al., 2017) with a max-attention layer (instead of softmax) is also homogeneous (however, practical models typically include residual connections, layer norms, softmax attention, and so on, that deviate from the above conditions). Yet, as we show empirically in Section 6.1, the conclusion of the above result holds to some extent.
Lemma 4 (Low-rank structure).
Let $\theta=\theta(\infty)$ be a stationary point of the $\lambda$-regularized loss function. For any $\lambda>0$, we have that
$\frac{1}{K}\sum_{k=1}^{K}\frac{\|W_{k}\|_{2}}{\|W_{k}\|_{F}}$ $\geq\left(\frac{\lambda}{L(\theta)^{1/2-1/K}}\right)^{1/2}\;.$ (2)
For exponential ($\ell(x)=e^{-x}$) and logistic ($\ell(x)=\ln(1+e^{-x})$) losses, we additionally have
$\frac{1}{K}\sum_{k=1}^{K}\frac{\|W_{k}\|_{2}}{\|W_{k}\|_{F}}$ $\geq\frac{\lambda e}{(\lambda e)^{1/K}e^{1/K}}\;.$ (3)
[Proof]
The above result implies that the average stable rank decreases as $\lambda$ increases. Timor et al. (2023) show low-rank structures for the global optimum of the minimum-rank interpolating solution assuming existence of a smaller “teacher” network with $K^{\prime}<K$ layers that interpolates data and Frobenius norm of its weight matrices are bounded by $C$. More specifically, they show that $\frac{1}{K}\sum_{k=1}^{K}\frac{\|W_{k}\|_{2}}{\|W_{k}\|_{F}}\geq\left(\frac{1}{C}\right)^{\frac{K^{\prime}}{K}}$. We can have an intuitive understanding of their result by noting that given a smaller interpolating network, the last layer of the network tends to be low-rank due to the Neural Collapse phenomena (Papyan et al., 2020). Then if we add more layers, the additional layers will remain low-rank as their activations will lie in a low-rank structure, and $\frac{1}{K}\sum_{k=1}^{K}\frac{\|W_{k}\|_{2}}{\|W_{k}\|_{F}}$ will approach one as we add more layers.
Results in Lemma 2 and Lemma 4 show a different phenomena. Remarkably, in both these results, irrespective of network capacity or performance, and as a consequence of weight decay, the weight matrices will become low-rank on average. Another interesting feature is that, unlike results of Timor et al. (2023), depth plays a minor role in establishing low-rank structures in Lemma 2 and Lemma 4. In Section 6.1, we present experiments that show that even if number of layers is small and network has a large number of classification mistakes, the average inverse stable-rank grows as weight decay parameter $\lambda$ increases.
5 Merging model parameters
Throughout this section, in addition to the training tuple $(x_{i},y_{i})_{i=1}^{n}$ we introduce $(x_{i}^{\prime},y_{i}^{\prime})_{i=1}^{n}$, and in addition to the regularized objective $L_{\lambda}$ we will introduce $L_{\lambda}^{\prime}$ defined with respect to the second training tuple.
As an instructive example first consider a linear prediction scenario where we are looking at predictors of a form $f_{\theta}(x)=\left\langle\theta,x\right\rangle$ and suppose that $\theta_{t}$ is obtained by GD minimizing $L_{\lambda}$, while $\theta_{t}^{\prime}$ is obtained by minimizing $L_{\lambda}^{\prime}$. Now given a test point $x$ that is sufficiently different from inputs $(x_{i}^{\prime})_{i=1}^{n}$, or assuming that $\max_{i}|\left\langle{x_{i}^{\prime}},x\right\rangle|\leq\varepsilon$ we have
$f_{\theta_{t}+\theta_{t}^{\prime}}(x)\approx f_{\theta_{t}}(x)$
and the same holds for some point $x^{\prime}$ sufficiently different from $(x_{i})_{i=1}^{n}$. This is derived in a straightforward way based on GD update rule
$\theta_{t+1}=(1-\lambda\eta)\theta_{t}-\eta\,\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime}(y_{i}\left\langle\theta_{t},x_{i}\right\rangle)x_{i}y_{i}$
while unrolling the recursive relationship we get
$\theta_{t}$ $=\theta_{0}(1-\eta\lambda)^{t}+\sum_{i=1}^{n}\alpha_{i}x_{i}\quad\text{where}\quad\alpha_{i}=\frac{1}{n}\sum_{s=0}^{t-1}\eta(1-\eta\lambda)^{t-s-1}\ell^{\prime}(y_{i}\left\langle\theta_{s},x_{i}\right\rangle)y_{i}~{}.$
The above identity tells us that the solution $\theta_{t}$ resides in the span of inputs. So, it is easy to see that prediction on the inputs from the other task is close to $\varepsilon$
$\left|\left\langle\theta_{t},x^{\prime}\right\rangle\right|$ $\leq\left|\left\langle\theta_{0},x^{\prime}\right\rangle\right|(1-\eta\lambda)^{t}+\varepsilon\,\frac{1-(1-\eta\lambda)^{t}}{\lambda}$
for a large enough $t$. In particular, the above implies that the gap of interest $|f_{\theta_{t}+\theta_{t}^{\prime}}(x)-f_{\theta_{t}}(x)|$ is also controlled by the upper bound in the display above. Note that a possibly large term $\left|\left\langle\theta_{0},x^{\prime}\right\rangle\right|$ is attenuated exponentially quickly by the weight decay, so its effect is negligible at the end of optimization. While in the linear case we could set $\theta_{0}=0$, in case of neural network learning setting initialization at 00 is atypical and therefore weight decay seems to have an important role in such scenarios. Another summand on the right hand side is $\varepsilon$-dependent and captures similarity between inputs from different tasks. If inputs are orthogonal this term disappears.
Next, we will demonstrate similar gaps for neural networks trained by GD and GF.
Shallow ReLU networks
Consider a shallow ReLU neural network
$f_{\theta}(x)=\left\langle u,(Wx)_{+}\right\rangle\qquad(x\in\mathbb{R}^{d})$
where hidden weight matrix $W\in\mathbb{R}^{m\times d}$ is a tunable parameters, and $u$ with $\|u\|\leq 1$ is fixed throughout training. In this scenario we obtain a merged predictor by simply summing hidden weight matrices of neural networks trained on different tasks. The intuition behind the argument in this case is that for an input $x$, the length $\|W_{t}^{\prime}x\|$ must be small because rows of $W_{t}^{\prime}$ lie in the span on $(x_{i}^{\prime})_{i}$ meanwhile each of these points is sufficiently different from $x$. We formalize this in the following lemma which applies to GD iterates:
Lemma 5.
Suppose that inputs $(x_{i})_{i}$ and $x^{\prime}$ are such that $\max_{i}\left|\left\langle x_{i},x^{\prime}\right\rangle\right|\leq\varepsilon$. Suppose that $W_{t}$ is a weight matrix of a shallow neural network obtained by running GD for $t$ steps given $(x_{i},y_{i})_{i=1}^{n}$. Then, for any $j\in[m],t\in\mathbb{N}$,
$\|W_{t}x^{\prime}\|\leq\|W_{0}x^{\prime}\|\,(1-\eta\lambda)^{t}+\varepsilon\,\frac{1-(1-\eta\lambda)^{t}}{\lambda}~{}.$
[Proof]
The lemma almost immediately implies
Corollary 1.
Suppose that inputs $(x_{i})_{i}$ and $(x_{i}^{\prime})_{i}$ satisfy $\max_{i,j}\left|\left\langle x_{i},x_{j}^{\prime}\right\rangle\right|\leq\varepsilon$ and that $W_{t}$ and $W_{t}^{\prime}$ are weight matrices of two shallow neural networks, obtained by running GD given datasets $(x_{i},y_{i})_{i}$ and $(x_{i}^{\prime},y_{i}^{\prime})_{i}$ respectively. Then, for any $x$ such that $\max\left|\left\langle x,x_{i}^{\prime}\right\rangle\right|\leq\varepsilon$ and any $t\geq 0$ we have
$|f_{\theta_{t}+\theta_{t}^{\prime}}(x)-f_{\theta_{t}}(x)|\leq\|W_{t}^{\prime}x\|\leq\|W_{0}^{\prime}x\|\,(1-\eta\lambda)^{t}+\varepsilon\,\frac{1-(1-\eta\lambda)^{t}}{\lambda}~{}.$
[Proof]
Deep linear networks
Next, we consider the predictor
$f_{\theta}(x)=\left\langle w_{K},W_{K-1}\cdots W_{1}x\right\rangle$
where each weight matrix is trained by GF dynamics as described in Eq. 1. To show the desired result we exploit a technical result of Arora et al. (2018) which translates GF dynamics for individual matrices into implicit GF dynamics for the end-to-end vector
$w(t)^{\top}=w_{K}(t)^{\top}W_{K-1}(t)\cdots W_{1}(t)~{}.$
The proof exploits the fact that $w(t)$ indeed lies in the span of inputs.
Theorem 1.
Assume that weight matrices are initialized such that for all $k\in[K-1]$
$W_{k+1}(0)^{\top}W_{k+1}(0)=W_{k}(0)W_{k}(0)^{\top}~{}.$
Then, for any $t\geq 0$ and any input $x$ such that $\max_{i}\left|\left\langle x_{i},x\right\rangle\right|\leq\varepsilon$ we have
$|f_{\theta(t)+\theta^{\prime}(t)}(x)-f_{\theta(t)}(x)|\leq|f_{\theta(0)}(x)|A_{1}\,e^{-\lambda Kt}+A_{2}\,\varepsilon$
where terms $A_{1},A_{2},B$ depend only on initialization, $\lambda$, and $K$ (see Theorem 3 for precise expressions). [Proof]
Theorem 1 gives the bound on the gap of the same form as in the shallow and linear case, that is an exponentially decaying term that arises because of the weight decay, and an $\varepsilon$-dependent term which captures similarity of inputs in different tasks. Unlike the previous cases, the theorem assumes a particular initialization, which is inherited from Arora et al. (2018): this assumptions is benign and is satisfied with high probability when entries of weight matrices are sampled from some symmetric distribution (e.g. isotropic Gaussian).
Deep ReLU networks
As our last result we show that parameter summation also works well in case of deep ReLU neural networks. Here, we consider a simpler scenario where such neural networks are trained by GF until convergence, that is to the zero-gradient of a regularized objective $L_{\lambda}$. [^3] Although we believe that it is possible to state a similar result for a finite time, working with a stationary point allows to greatly simplify the proofs. In other words we look at the stationary points of $L_{\lambda}$. Note that convergence to the local minimum is not required in our analysis.
However, we require two fairly mild technical assumptions. The first, 1 says that on-average margins are positive, and we will assume that this is satisfied at some time $t_{0}$. Note that this implicitly assumes that the capacity of the predictor has to be large enough to correctly classify at least a fraction of training points. This type of assumption is common in the literature, for instance (Lyu and Li, 2020, Ji and Telgarsky, 2020) assume that $L(\theta(0))\leq\ell(0)$.
The second assumption, 2, captures scenarios where the set of positive activations converges to a non-empty set. This prevents degenerate cases where all activations eventually vanish (which is possible if weight decay is too strong).
Assumption 1 (On-average positive margin).
We say that $\theta$ satisfies on-average positive margin condition on training tuple $(x_{1},y_{i}),\ldots,(x_{n},y_{n})$ when
$-\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime}(y_{i}\,f_{\theta}(x_{i}))\,y_{i}\,f_{\theta}(x_{i})\geq 0~{}.$
Definition 1 (Smallest positive normalized activation).
The smallest positive normalized activation of ReLU neural network at layer $k$, provided input $x$, is defined as
$\gamma_{k}(x;\theta)=\min\left\{\frac{\left\langle e_{j},W_{k}h_{k-1}(x;\theta)\right\rangle}{\|h_{k-1}(x;\theta)\|}~{}:~{}\left\langle e_{j},W_{k}h_{k-1}(x;\theta)\right\rangle>0~{},j\in[m]\right\}~{}.$
Assumption 2 (Convergence of positive activations).
We say that positive activations at layer $k$ converge given input $x$, if $\lim_{t\to\infty}\gamma_{k}(x;\theta(t))=c$ for $c>0$.
The following is the main result of this section:
Theorem 2.
Suppose that parameters of two ReLU neural networks $\theta,\theta^{\prime}$ are limiting solutions of GF minimizing $L_{\lambda}$ and $L^{\prime}_{\lambda}$ respectively.
  • Assume that there exists time $t_{0}$ such that both $\theta(t_{0}),\theta^{\prime}(t_{0})$ satisfy on-average positive margin condition of 1.
  • Assume that activations at layer $k$ converge in a sense of 2 for both networks given inputs $x,x^{\prime}$ respectively.
  • Finally, assume that inputs $x,x^{\prime}$ satisfy $\left\langle x_{i},x^{\prime}\right\rangle=0$ and $\left\langle x_{i}^{\prime},x\right\rangle=0$.
Then, there exists a tuning of weight decay parameters $(\lambda_{t})_{t}$ such that $\lambda_{t}\to 0$ as $t\to 0$ such that $h_{k}(x;\,\theta+\theta^{\prime})=h_{k}(x;\,\theta)$ and similarly for $(x^{\prime},\theta^{\prime})$. [Proof]
More precisely, the tuning of $\lambda_{t}$ mentioned in the theorem is of order $\lambda_{t}=\Omega(\frac{1}{\gamma_{k}\sqrt{t}})$ where $\gamma_{k}<\min_{t\geq t_{0}}\gamma_{k}(x,\theta(t))$, $\min_{t\geq t_{0}}\gamma_{k}(x,\theta^{\prime}(t))$.
This result is somewhat stronger than the previous ones as it shows that merging is effective at any layer, and since we consider stationary points, we achieve an identity.
The proof of Theorem 2 is conceptually similar to that of the shallow neural network case (Lemma 5). In particular, in the shallow case we were relying on the fact that $\|W_{k}x^{\prime}\|$ is small, meanwhile here we need to show that $\|W_{k}h_{k-1}(x^{\prime};\theta^{\prime})\|=0$. Alignment (Lemma 1) tells us that rows of $W_{k}$ will lie in the span of activation vectors $h_{k-1}(x_{i};\theta)$, and so the missing piece is to show that $h_{k-1}(x_{i};\theta)$ is orthogonal to $h_{k-1}(x^{\prime};\theta^{\prime})$. This is largely the consequence of the weight decay: In Lemma 8 we establish that dot product between activations vanishes exponentially quickly under appropriate conditions (1 and 2).
6 Experiments
6.1 Norm and rank structure
In this section we aim to investigate whether some of the biases discussed in Section 4 extend do very deep neural networks. In particular, we aim to verify whether norm preservation and low stable-ranks can be found in some (smaller) LLMs. Even though assumptions of Section 4 might not be satisfied, we still find that in several LLMs we observe low stable-ranks (drastically smaller than the dimension of the matrix), while norm preservation appears in most of the layers (MLP and attention ones). In these experiments we look at publicly available pre-trained BERT with $\approx$ 110M parameters, GPT2-Large with $\approx$ 774M , GPT2-XL with $\approx$ 1.5B , RoBERTa with $\approx$ 125M , Phi-2 with $\approx$ 2.8B , GPT-J with $\approx$ 6B (we used Hugging Face “Transformers” library (Wolf et al., 2020)).
Each of these ‘transformer’ models consists of a sequence of a so-called encoder layers, where each encoder layer consists of a self-attention layer followed by a fully-connected neural network. Self-attention layer is given by a matrix-to-matrix function $Q\mapsto\text{softmax}\big{(}QK^{T}/\sqrt{\text{columns}(K)}\big{)}V$ where $(Q,V)$ are parameter matrices and softmax is taken row-wise. Each self-attention layer is followed by a feed-forward fully connected neural network consisting of two linear layers with a non-linear activation function (e.g., ReLU or a smooth activation function) in between. In the context of transformer architecture, $Q,K,V$ are known as query, key, and value matrices. In Appendix C we provide a table Table 1 that summarizes architectural details of these models.
While some of the results are given in Appendix C, here in Figure 1 we provide results for two largest neural networks. In these plots, Q, K, V, O, and QKV denote Query, Key, Value, Output, and the concatenation of Query, Key, Value matrices of the attention layer, respectively. Weight matrices of MLP layers are denoted by M1 and M2.
The plots show that the Frobenius norms of QKV, M1, and M2 matrices are generally in the same order. Although this might appear puzzling at first, it is in fact consistent with Lemma 3: the lemma is a statement about norms of layer weights when the output of the network can be written as a product of those layers. In order to apply the lemma to a Transformer architecture, we should consider the QKV matrix as a layer weight instead of considering the individual attention matrices. All weight matrices generally have small stable-ranks, although the values can be different for different parameter type and layer.
There are two observations for which we currently have no explanation: (1) The Frobenius norms of the same parameter type (even Q, K, V matrices) remain largely unchanged across layers. (2) Even though Lemma 3 suggests the O matrix should have similar Frobenius norms as QKV, M1, and M2 matrices, the plots show quite different values for norms of O matrices.
Figure 1: Stable ranks and Frobenius norms of different weight matrices in pretrained Phi-2 (first row) and GPT-J (second row) models.
[…]
Figure 2: Low rank induced by weight decay.
6.2 Model merging
In our experiments we aim to verify four hypotheses:
  • Training two neural networks on different tasks (with nearly orthogonal inputs), and summing the weights results in a combined neural network that performs nearly as well on each of the tasks. Here the performance is understood in terms of the training loss.
  • In a contrast, training two neural networks on the same task and performing the merging as discussed above leads to the combined predictor that performs poorly on both tasks.
  • This is behavior is enabled by the weight decay.
  • Using weight decay leads to a low stable rank of weight matrices.
Data.
[…]
Model and training.
In all the experiments Fully-connected ReLU neural network with inputs $d=100$, two hidden layers of sizes $(1000,100)$, and a scalar output.
In each experiment on synthetic data, models are trained by GD with step size $\eta=1$ over $10^{5}$ steps. In all the experiments excepts the one where weight decay varies, weight decay parameter $\lambda$ is set to $10^{-4}$. On Fashion MNIST dataset, step size is set as $\eta=0.1$ while $\lambda=10^{-3}$. All experiments are repeated on $10$ random draws of the sample, and we report standard deviations in all plots.
Figure 3: First row: Training neural networks on different tasks (orthogonal inputs) (left) vs. the same task (right) and merging the parameters by adding weight matrices. The resulting network performs well on different tasks after sufficiently many iterations, while given the same task, it does not. Second row: this effect manifests when weight decay strength is sufficiently large (left). Stable rank of each weight matrix converges to a small value. Merged network matches the stable rank of individual networks (right).
Discussion.
On both datasets, we consistently observe that summing weight matrices originating from different tasks, enables small logistic loss for merged models trained on all tasks after sufficiently long training (first row in Figure 3). This seems to be in part enabled by orthogonality of inputs since, in contrast, when inputs are not orthogonal and labels (or labeling functions) are different, merged models do not get close in performance to distinct task-specific models. Another component that enables this gap is weight decay: In the second row of Figure 3 we observe that when weight decay strength is not sufficient, the gap between losses of merged and original models is substantial. Finally, we observe that the stable rank converges to a value much smaller than the actual rank at the stage when performance of merged model actually gets close to the performance of original model. This is another observation in support of our hypothesis that low-rank bias is crucial for model merging.
7 Additional related work
Weight averaging
Several prior works have considered building a final model by averaging weights of different models (Utans, 1996, Pavel Izmailov and Wilson, 2018, Alexandre Rame and Cord, 2022, Mitchell Wortsman and Schmidt, 2022, Zafir Stojanovski and Akata, 2022, Gabriel Ilharco and Schmidt, 2022, Gabriel Ilharco and Farhadi, 2023, Chung et al., 2023, Douillard et al., , Alexandre Rame and Cord, 2023, Ram’e et al., 2024, Ramé et al., 2024). To the best of our knowledge, prior works mostly consider settings where several models are trained on data from the same or similar tasks. Their main idea is that the randomization in data or parameter initialization leads to perturbations in the learned weights, and taking their average leads to more stable solutions. This is also known as “linear mode connectivity" (Frankle et al., 2020, Neyshabur et al., 2020). In contrast, we consider models trained on entirely different tasks, and we take the weights’ summation instead of their average. Our explanation for the success of model merging is also very different and relies on the low-rank structures of the learned weights.
Neural Collapse
Low-rank bias is intimately related to a so-called Neural Collapse (NC), which is a form of clustering within the feature space of a neural network (Papyan et al., 2020). NC hypothesis suggests that the network refines its representations so that inputs belonging to the same class (in context of classification) are pulled closer together, forming distinct clusters. Simultaneously, the network pushes the cluster centers (class means) away from each other, maximizing their separation.
Some of the existing results are mostly in the Unconstrained Feature Model (UFM) where only the last-layer features and the output of linear classifier are learnable. This is in fact equivalent to learning a shallow linear network. Most studies show that the global optima of the loss function satisfies the conditions of neural collapse, without studying the dynamics of gradient descent.
Yang et al. (2022) demonstrated that in the Universal Feature Manifold (UFM) setting, fixing the linear output layer to satisfy a specific condition induces neural collapse in the feature vectors that minimize the loss, even with imbalanced data. This fixed output layer approach allows them to extend typical neural collapse results, which often assume balanced data, to the more challenging imbalanced case.
Rangamani and Banburski-Fahey (2022) showed that for ReLU deep networks trained on balanced datasets using gradient flow with weight decay and a squared loss, critical points satisfying a "Symmetric Quasi-interpolation" assumption also exhibit neural collapse. This assumption, which posits the existence of a classifier whose output depends only on the class label and not the specific data point, is presented as a key condition for their result.
Notably, model merging discussed here differs from weight averaging (Utans, 1996): weight averaging relies on learned weights converging to vicinity of each other, so taking their average leads to a more stable solution. In our case, learned weights are far from each other, and the average weights often do not perform well in practice. As our theory suggests, we should sum the weights up, which in fact leads to good performance.
8 Conclusions
In this work we examined the role of $L2$ regularization in training of deep neural networks with logistic loss. We investigated a surprising phenomenon: merging two neural networks trained on sufficiently different tasks by simply adding their respective weight matrices results in a predictor that performs well on both tasks simultaneously. As we attributed the explanation of this to the low-rank bias arising in weight matrices, we also established that $L2$ regularization leads to weight matrices of a low stable rank.
These observations open up some interesting possibilities, especially in multitask learning with large models, such as large language models, and distributed optimization. We believe our proof technique for deep ReLU neural networks ( Theorem 2) can be refined, potentially removing some of the assumptions, such as 2.
Appendix A Proofs from Section 4
Proof of Lemma 1.
If $L_{\lambda}$ is differentiable and $H$-smooth with $H>0$, by the descent lemma,
$\frac{1}{2H}\|\nabla L_{\lambda}(\theta_{t})\|^{2}\leq L_{\lambda}(\theta_{t})-L_{\lambda}(\theta_{t+1})\;.$
By summing over all terms and the fact that losses are non-negative,
$\frac{1}{2H}\sum_{t=0}^{\infty}\|\nabla L_{\lambda}(\theta_{t})\|^{2}\leq L_{\lambda}(\theta_{0})\;.$
Boundedness of the above sum implies that $\|\nabla L_{\lambda}(\theta_{t})\|\rightarrow 0$ as $t\rightarrow\infty$. Therefore, at the limit we have the alignment of the weight and the gradient vectors, $\lambda\theta=-\nabla L(\theta)$. ∎
Proof of Lemma 2.
Let $w^{(i)\top}=w_{K}^{\top}W_{K-1}\dots W_{i}$. Therefore, $f_{\theta}(x)=\left\langle w^{(i)},W_{i-1}\dots W_{1}x\right\rangle$. Let $r_{i}(\theta):=-y_{i}\ell^{\prime}(y_{i}f_{\theta}(x_{i})/(n\lambda)$. By alignment, we have
$W_{1}$ $=w^{(2)}\sum_{i=1}^{n}r_{i}(\theta)x_{i}^{\top}\,,$
$W_{2}$ $=w^{(3)}\sum_{i=1}^{n}r_{i}(\theta)(W_{1}x_{i})^{\top}\,,$
$\vdots$
$W_{K-1}$ $=w^{(K)}\sum_{i=1}^{n}r_{i}(\theta)(W_{K-2}\dots W_{1}x_{i})^{\top}\,,$
$w_{K}$ $=\sum_{i=1}^{n}r_{i}(\theta)(W_{K-1}\dots W_{1}x_{i})^{\top}\;.$
It’s easy to see that all weight matrices above are rank-1. ∎
Proof of Lemma 3.
Lemma 1 holds also for the parameters of layer $k$: $\lambda W_{k}=-\nabla_{W_{k}}L(\theta)$. Therefore,
$\lambda\|W_{k}\|_{F}^{2}$ $=\lambda\,\operatorname*{tr}(W_{k}W_{k}^{\top})$
$=-\operatorname*{tr}(\nabla_{W_{k}}L(\theta)W_{k}^{\top})$
$=-\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime}(y_{i}f_{\theta}(x_{i}))y_{i}\operatorname*{tr}(\nabla_{W_{k}}f_{\theta}(x_{i})W_{k}^{\top})$
$=-\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime}(y_{i}f_{\theta}(x_{i}))y_{i}f_{\theta}(x_{i})\,,$
where the last step holds by the fact that for $f_{\theta}$ locally Lipschitz and positively homogeneous, $f_{\theta}(x)=\operatorname*{tr}(\nabla_{W_{k}}f_{\theta}(x_{i})W_{k}^{\top})$ (see e.g. Lemma 9.2 of Telgarsky (2021)). Given that the equation holds independently of layer index $k$, weight matrices all have the same Frobenius norm. ∎
Proof of Lemma 4.
By AM-GM inequality,
$\lambda\|W_{k^{\prime}}\|_{F}^{2}\leq\prod_{k=1}^{K}\|W_{k}\|_{2}\leq\left(\frac{1}{K}\sum_{k=1}^{K}\|W_{k}\|_{2}\right)^{K}\;.$
Therefore,
$\frac{1}{K}\sum_{k=1}^{K}\frac{\|W_{k}\|_{2}}{\|W_{k}\|_{F}}$ $=\frac{1}{B}\cdot\frac{1}{K}\sum_{k=1}^{K}\|W_{k}\|_{2}$
$\geq\frac{1}{B}\left(\lambda B^{2}\right)^{1/K}\;.$
To get Equation 2, we note that by Lemma 3,
$\lambda\|W_{k}\|_{F}^{2}$ $=-\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime}(y_{i}f_{\theta}(x_{i}))y_{i}f_{\theta}(x_{i})$
$\leq\frac{1}{n}\sum_{i=1}^{n}\sqrt{\ell(y_{i}f_{\theta}(x_{i}))}$
$\leq\sqrt{\frac{1}{n}\sum_{i=1}^{n}\ell(y_{i}f_{\theta}(x_{i}))}$
$=\sqrt{L(\theta)}\;.$
To get Equation 3, we again use Lemma 3 and get that for the exponential and logistic losses, we have $B\leq 1/(e\lambda)$.
Appendix B Proofs from Section 5
B.1 Proof of Lemma 5
Observe that GD update rule is
$W_{s+1}=(1-\eta\lambda)W_{s}-\eta\,\frac{1}{n}\sum_{i=1}^{n}y_{i}\,\ell_{i}(y_{i}\,f_{\theta_{s}}(x_{i}))\,D_{s,i}\,u\,x_{i}^{\top}$
where $D_{s,i}=\operatorname*{diag}(\mathbbm{1}(W_{s}x_{i}>0))$. Then
$W_{s+1}x^{\prime}=(1-\eta\lambda)W_{s}x^{\prime}-\eta\,\frac{1}{n}\sum_{i=1}^{n}y_{i}\,\ell_{i}(y_{i}\,f_{\theta_{s}}(x_{i}))D_{s,i}\,u\left\langle x_{i},x^{\prime}\right\rangle$
which together with Cauchy-Schwartz inequality implies
$\|W_{s+1}x^{\prime}\|$ $\leq(1-\eta\lambda)\|W_{s}x^{\prime}\|+\eta\,\frac{1}{n}\sum_{i=1}^{n}|y_{i}\,\ell_{i}(y_{i}\,f_{\theta_{s}}(x_{i}))|\|D_{s,i}\|_{2}\,\|u\|\left|\left\langle x_{i},x^{\prime}\right\rangle\right|$
$\leq(1-\eta\lambda)\|W_{s}x^{\prime}\|+\varepsilon\,\eta~{}.$
Observe that relation $x_{s+1}\leq a_{s}x_{s}+b_{s}$ unwinds from $t$ to $t_{0}$ as $x_{t}\leq x_{t_{0}}\prod_{k=t_{0}}^{t-1}a_{k}+\sum_{s=t_{0}}^{t-1}b_{s}\prod_{k=s+1}^{t-1}a_{k}$. So,
$\|W_{t}x^{\prime}\|$ $\leq\|W_{0}x^{\prime}\|\,(1-\eta\lambda)^{t}+\varepsilon\,\eta\,\sum_{s=0}^{t-1}\prod_{k=s+1}^{t-1}(1-\eta\lambda)$
$=\|W_{0}x^{\prime}\|\,(1-\eta\lambda)^{t}+\varepsilon\,\eta\,\sum_{s=0}^{t-1}(1-\eta\lambda)^{t-s-1}~{}.$
B.2 Proof of Corollary 1
First observe that
$|f_{\theta_{t}+\theta_{t}^{\prime}}(x)-f_{\theta_{t}}(x)|$ $\leq\left|\left\langle u,(W_{t}x+W_{t}^{\prime}x)_{+}\right\rangle-\left\langle u,(W_{t}x)_{+}\right\rangle\right|$
$\leq\|(W_{t}x+W_{t}^{\prime}x)_{+}-(W_{t}x)_{+}\|$ (Cauchy-Schwartz inequality)
$\leq\|W_{t}^{\prime}x\|$
where the last inequality comes by a basic fact about ReLUs (see Section 2). Now, since $\ell$ is $1$-Lipschitz,
$\ell\left(yf_{\theta_{t}+\theta_{t}^{\prime}}(x)\right)-\ell\left(yf_{\theta_{t}}(x)\right)$ $=\ell\left(y\left\langle u,(W_{t}x+W_{t}^{\prime}x)_{+}\right\rangle\right)-\ell\left(y\left\langle u,(W_{t}x)_{+}\right\rangle\right)$
$\leq|f_{\theta_{t}+\theta_{t}^{\prime}}(x)-f_{\theta_{t}}(x)|$
$\leq\|W_{t}^{\prime}x\|~{}.$
At this point, we apply losses over $(x_{i},y_{i})_{i}$ and average to have
$L(\theta_{t}+\theta_{t}^{\prime})\leq L(\theta_{t})+\frac{1}{n}\sum_{i=1}^{n}\|W_{t}^{\prime}x_{i}\|~{}.$
Now we use Lemma 5 to control $\|W_{t}^{\prime}x_{i}\|$. ∎
B.3 Proof of Theorem 1
Theorem 1 is a direct corollary of the following theorem which we show in this section:
Theorem 3.
Assume that weight matrices are initialized such that for all $k\in[K-1]$
$W_{k+1}(0)^{\top}W_{k+1}(0)=W_{k}(0)W_{k}(0)^{\top}~{}.$
Let $\ell$ be a logistic loss and let $L(\theta(0))\leq C$. Then, for any $t\geq 0$ and any input $x^{\prime}$ such that $\max_{i}\left|\left\langle x_{i},x^{\prime}\right\rangle\right|\leq\varepsilon$ we have
$|\left\langle w(t),x^{\prime}\right\rangle|\leq|\left\langle w(0),x^{\prime}\right\rangle|A_{1}\,e^{-\lambda Kt}+A_{2}\,\varepsilon$
where
$A_{1}$ $=\exp\left(\frac{B^{2-\frac{2}{K}}(K-1)(1+\lambda K)C}{2\lambda K}\right)~{},$
$A_{2}$ $=2B^{2-\frac{2}{K}}C\,A_{1}\left(\frac{1-e^{-\lambda Kt/2}}{\lambda K}\right)~{},$
$B$ $=\|w(0)\|^{2}+\frac{C}{\lambda}~{}.$
[Proof]
In the following for the end-to-end vector $w(t)^{\top}=w_{K}(t)^{\top}W_{K-1}(t)\cdots W_{1}(t)$ we use notation $L^{1}$ to denote its loss:
$L^{1}(w(t))=\frac{1}{n}\sum_{i=1}^{n}\ell(y_{i}\left\langle w(t),x_{i}\right\rangle)~{}.$
Note that $L^{1}(w(t))=L(\theta(t))$. Proof relies on the following crucial result connecting per-layer updates to updates of the product matrix:
Theorem 4 (Arora et al. (2018, Theorem 1)).
Assume that weight matrices are initialized (at time $t_{0}$) in such a way that they satisfy for all $k\in[K-1]$.
$W_{k+1}(t_{0})^{\top}W_{k+1}(t_{0})=W_{k}(t_{0})W_{k}(t_{0})^{\top}$
Then, under dynamics with updates as in Eq. 1 for the end-to-end matrix we have
$\dot{W}(t)=-\lambda K\cdot W(t)-\sum_{k=1}^{K}\left(W(t)W(t)^{\top}\right)^{\frac{k-1}{K}}\cdot\nabla L^{1}(W)\cdot\left(W(t)^{\top}W(t)\right)^{\frac{K-k}{K}}~{}.$
B.3.1 Proof of Theorem 3
We start from adapting Theorem 4 to our case to get $w(t)$ and then get an identity for $\left\langle w(t),x^{\prime}\right\rangle$ by solving the resulting differential equation. Once we get dependence on $\varepsilon$ we must ensure that the remaining terms (arising from the solution to differential equation) are non-divergent, which we will do through the stationary point convergence argument.
Theorem 4 gives us
$\dot{w}(t)^{\top}$ $=-\lambda K\cdot w(t)^{\top}-\|w(t)\|^{\frac{2(K-1)}{K}}\cdot\nabla L^{1}(w(t))^{\top}$
$\qquad-\sum_{k=1}^{K-1}\|w(t)\|^{\frac{2(k-1)}{K}}\cdot\nabla L^{1}(w(t))^{\top}\left(w(t)w(t)^{\top}\right)^{\frac{K-k}{K}}$
$=-\lambda K\cdot w(t)^{\top}-\|w(t)\|^{2-\frac{2}{K}}\cdot\nabla L^{1}(w(t))^{\top}$
$\qquad-\sum_{k=1}^{K-1}\|w(t)\|^{\frac{2(k-1)}{K}+\frac{2(K-k)}{K}}\cdot\nabla L^{1}(w(t))^{\top}\left(\frac{w(t)w(t)^{\top}}{\|w(t)\|^{2}}\right)^{\frac{K-k}{K}}$
$=-\lambda K\cdot w(t)^{\top}-\|w(t)\|^{2-\frac{2}{K}}\cdot\nabla L^{1}(w(t))^{\top}$
$\qquad-(K-1)\|w(t)\|^{2-\frac{2}{K}}\cdot\left\langle\nabla L^{1}(w(t)),w(t)\right\rangle w(t)^{\top}~{}.$
Put another way,
$\dot{w}(t)=\left(-\lambda K-(K-1)\|w(t)\|^{2-\frac{2}{K}}\,\left\langle\nabla L^{1}(w(t)),w(t)\right\rangle\right)\,w(t)-\|w(t)\|^{2-\frac{2}{K}}\,\nabla L^{1}(w(t))~{}.$ (4)
Taking dot product with $x^{\prime}$ gives
$\left\langle\dot{w}(t),x^{\prime}\right\rangle=a(t)\,\left\langle w(t),x^{\prime}\right\rangle+b(t)$ (5)
where we introduce abbreviations
$a(t)$ $:=-\lambda K-(K-1)\|w(t)\|^{2-\frac{2}{K}}\left\langle\nabla L^{1}(w(t)),w(t)\right\rangle$
$b(t)$ $:=-\|w(t)\|^{2-\frac{2}{K}}\left\langle\nabla L^{1}(w(t)),x^{\prime}\right\rangle~{}.$
Solving Eq. 5 we get
$\left\langle w(t),x^{\prime}\right\rangle$ $=\left\langle w(t_{0}),x^{\prime}\right\rangle\,e^{\int_{t_{0}}^{t}a(s)\mathop{}\!\mathrm{d}s}+\int_{t_{0}}^{t}b(s)\,e^{\int_{s}^{t}a(r)\mathop{}\!\mathrm{d}r}\mathop{}\!\mathrm{d}s~{}.$
Assuming for a moment that $\|w(t)\|\leq B$ (where $B$ will be determined later) Cauchy-Schwartz inequality gives
$a(t)\leq-\lambda K+(K-1)B^{1-\frac{1}{K}}\|w(t)\|^{1-\frac{1}{K}}|\left\langle\nabla L^{1}(w(t)),w(t)\right\rangle|~{}.$
On the other hand, using the fact that for logistic loss function $|\ell^{\prime}(z)|\leq\ell(z)$,
$b(t)$ $=\|w(t)\|^{2-\frac{2}{K}}\,\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime}(y_{i}\left\langle w(t),x_{i}\right\rangle)y_{i}\left\langle x_{i},x^{\prime}\right\rangle$
$\leq\varepsilon\cdot B^{2-\frac{2}{K}}\,\frac{1}{n}\sum_{i=1}^{n}\ell(y_{i}\left\langle w(t),x_{i}\right\rangle)$
$=\varepsilon\cdot B^{2-\frac{2}{K}}\,L^{1}(w(t))~{}.$
So, it is left to show that the term $\int_{t_{0}}^{t}a(s)\mathop{}\!\mathrm{d}s$ does not diverge. To this end, we show the following (with proof at the end of the section):
Lemma 6 (Stationary point convergence for dynamics in Eq. 4).
$(K-1)\int_{s}^{t}\|w(r)\|^{1-\frac{1}{K}}|\left\langle\nabla L^{1}(w(r)),w(r)\right\rangle|\mathop{}\!\mathrm{d}r\leq\sqrt{(K-1)(1+\lambda K)L^{1}(w(0))\cdot(t-s)}~{}.$
Using this lemma to bound $\int a(s)\mathop{}\!\mathrm{d}s$ gives
$|\left\langle w(t),x^{\prime}\right\rangle|$ $\leq|\left\langle w(0),x^{\prime}\right\rangle|\,\underbrace{\exp\left(-\lambda K\cdot t+B^{1-\frac{1}{K}}\sqrt{(K-1)(1+\lambda K)C\cdot t}\right)}_{(i)}$
$\qquad+\varepsilon\cdot B^{2-\frac{2}{K}}C\,\underbrace{\int_{0}^{t}\exp\left(-\lambda K\cdot(t-s)+B^{1-\frac{1}{K}}\sqrt{(K-1)(1+\lambda K)C\cdot(t-s)}\right)\mathop{}\!\mathrm{d}s}_{(ii)}$
where we also assumed that $\sup_{t}L^{1}(w(t))\leq C$. At this point we will bound $(i)$ and $(ii)$ by using the fact that
$e^{-at+b\sqrt{t}}\leq e^{\frac{b^{2}}{2a}}\,e^{-at/2}\qquad(a,b,t>0)~{}.$
that comes from (choosing $p$ such that $\frac{b}{2\sqrt{p}}=a/2$):
Proposition 1.
For any $a,b,t,p>0$,
$-at+b\sqrt{t}\leq\left(-a+\frac{b}{2\sqrt{p}}\right)t+\frac{b}{2}\sqrt{p}~{}.$
In particular this gives
$(ii)\leq 2\left(\frac{1-e^{-\lambda Kt/2}}{\lambda K}\right)\exp\left(\frac{B^{2-\frac{2}{K}}(K-1)(1+\lambda K)C}{2\lambda K}\right)$
As promised, the final bit is to give $B$. Since objective is non-increasing
$\lambda\|w(t)\|^{2}\leq\lambda\|w(t)\|^{2}+L^{1}(w(t))\leq\lambda\|w(0)\|^{2}+L^{1}(w(0))\qquad\implies\qquad B=\|w(0)\|^{2}+\frac{L^{1}(w(0))}{\lambda}~{}.$
Proof of Lemma 6..
Using the chain rule together with Eq. 4 we have
$\frac{\mathop{}\!\mathrm{d}L^{1}(w(t))}{\mathop{}\!\mathrm{d}t}$ $=\left\langle\nabla L^{1}(w(t)),\dot{w}(t)\right\rangle$
$=-\lambda K\left\langle\nabla L^{1}(w(t)),\dot{w}(t)\right\rangle-(K-1)\|w(t)\|^{2-\frac{2}{K}}\,\left\langle\nabla L^{1}(w(t)),w(t)\right\rangle^{2}$
$-\|w(t)\|^{2-\frac{2}{K}}\,\|\nabla L^{1}(w(t))\|^{2}$
and so
$L^{1}(w(t))-L^{1}(w(0))$ $=-\lambda K\int_{0}^{t}\left\langle\nabla L^{1}(w(s)),\dot{w}(s)\right\rangle\mathop{}\!\mathrm{d}s$
$-(K-1)\int_{0}^{t}\|w(s)\|^{2-\frac{2}{K}}\,\left\langle\nabla L^{1}(w(s)),w(s)\right\rangle^{2}\mathop{}\!\mathrm{d}s$
$-\int_{0}^{t}\|w(s)\|^{2-\frac{2}{K}}\,\|\nabla L^{1}(w(s))\|^{2}\mathop{}\!\mathrm{d}s~{}.$
Note also that the chain rule gives $\int_{0}^{t}\left\langle\nabla L^{1}(w(s)),\dot{w}(s)\right\rangle\mathop{}\!\mathrm{d}s=L^{1}(w(t))-L^{1}(w(0))$ which gives
$(K-1)\int_{0}^{t}\|w(s)\|^{2-\frac{2}{K}}\,\left\langle\nabla L^{1}(w(s)),w(s)\right\rangle^{2}\mathop{}\!\mathrm{d}s+\int_{0}^{t}\|w(s)\|^{2-\frac{2}{K}}\,\|\nabla L^{1}(w(s))\|^{2}\mathop{}\!\mathrm{d}s$
$\qquad=(1+\lambda K)\left(L^{1}(w(0))-L^{1}(w(t))\right)~{}.$
Applying Jensen’s inequality completes the proof. ∎
B.4 Proof of Theorem 2
Before proving the theorem we will require few basic tools about GF. The first tool used in the proof is stationarity of GF and its consequences:
Proposition 2.
For $W(t)$ obtained through dynamics $\dot{W}(t)=-\nabla F(W(t))$, where $F$ is uniformly continuous possibly non-differentiable and satisfies assumptions of Clarke differentiation (see Section 2), for almost every $t\geq 0$ we have
$\int_{0}^{t}\|\nabla F(W(s))\|_{F}^{2}=F(W(0))-F(W(t))~{}.$
In particular this implies that
  • $t\mapsto F(W(t))$ is non-increasing.
  • If $F(W(0))<\infty$ we have $\|\nabla F(W(t))\|_{F}\to 0$ as $t\to\infty$.
Proof.
The first statement is
$F(W(t))-F(W(0))=-\int_{0}^{t}\min\{\|g\|^{2}~{}:~{}g\in\partial F(W(s))\}\mathop{}\!\mathrm{d}s=-\int_{0}^{t}\|\nabla F(W(s))\|^{2}\mathop{}\!\mathrm{d}s~{}.$
The second statement is immediate as $F(W(0))$ is bounded. ∎
The above implies a stationarity result for GF with weight decay:
Lemma 7.
For $W_{k}(s)$ obtained through dynamics of Eq. 1, for any $k\in[K]$ any $t\geq t_{0}$,
$\int_{t_{0}}^{t}\|\nabla_{W_{k}}L(\theta(s))\|^{2}\mathop{}\!\mathrm{d}s+\lambda\int_{t_{0}}^{t}\left\langle\nabla_{W_{k}}L(\theta(s)),W_{k}(s)\right\rangle\mathop{}\!\mathrm{d}s=L(\theta(t_{0}))-L(\theta(t))~{}.$
Moreover, assuming that $\theta\mapsto f_{\theta}$ is homogeneous and satisfies 1 at initial time $t_{0}$, we have
$\int_{t_{0}}^{t}\|\nabla_{W_{k}}L(\theta(s))\|^{2}\mathop{}\!\mathrm{d}s\leq L(\theta(t_{0}))-L(\theta(t))~{}.$
Proof.
From Proposition 2 we have
$\int_{t_{0}}^{t}\|\nabla_{W_{k}}L(\theta(s))+\lambda W_{k}(s)\|^{2}\mathop{}\!\mathrm{d}s=C:=L(\theta(t_{0}))+\frac{\lambda}{2}\|W_{k}(t_{0})\|^{2}-L(\theta(t))-\frac{\lambda}{2}\|W_{k}(t)\|^{2}$
or
$\int_{t_{0}}^{t}\|\nabla_{W_{k}}L(\theta(s))\|^{2}\mathop{}\!\mathrm{d}s+2\lambda\int_{t_{0}}^{t}\left\langle\nabla_{W_{k}}L(\theta(s)),W_{k}(s)\right\rangle\mathop{}\!\mathrm{d}s+\lambda^{2}\int_{t_{0}}^{t}\|W_{k}(s)\|^{2}\mathop{}\!\mathrm{d}s=C$
On the other hand,
$\frac{1}{2}\|W_{k}(t)\|^{2}-\frac{1}{2}\|W_{k}(t_{0})\|^{2}$ $=\frac{1}{2}\int_{t_{0}}^{t}\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}s}\|W_{k}(s)\|^{2}\mathop{}\!\mathrm{d}s$
$=\int_{t_{0}}^{t}\left\langle W_{k}(s),\dot{W}_{k}(s)\right\rangle\mathop{}\!\mathrm{d}s$
$=-\lambda\int_{t_{0}}^{t}\|W_{k}(s)\|^{2}\mathop{}\!\mathrm{d}s-\int_{t_{0}}^{t}\left\langle W_{k}(s),\nabla_{W_{k}}L(\theta(s))\right\rangle\mathop{}\!\mathrm{d}s$
which implies the first statement.
The second statement comes by observing that by homogeneity of $f$,
$\left\langle\nabla_{W_{k}}L(\theta(s)),W_{k}(s)\right\rangle$ $=-\frac{1}{n}\sum_{i=1}^{n}y_{i}\ell^{\prime}(y_{i}f_{\theta(s)}(x_{i}))\left\langle\nabla_{W_{k}}f_{\theta(s)}(x_{i}),W(s)\right\rangle$
$=-\frac{1}{n}\sum_{i=1}^{n}y_{i}\ell^{\prime}(y_{i}f_{\theta(s)}(x_{i}))f_{\theta(s)}(x_{i})\geq 0$
by assumption of the lemma. ∎
The second crucial component is to show that activations of two different neural networks trained on different tasks are approximately orthogonal.
Lemma 8.
Suppose that parameters for two ReLU neural networks $\theta(t),\theta^{\prime}(t)$ are obtained by running GF given training tuples $(x_{i},y_{i})_{i=1}^{n}$ and $(x_{i}^{\prime},y_{i}^{\prime})_{i=1}^{n}$ respectively. Suppose that at time $t_{0}$ both $\theta(t),\theta^{\prime}(t)$ satisfy on-average positive margin condition of 1. Then, for any two inputs $x,x^{\prime}$, assuming the smallest positive normalized activation of both networks satisfy (see Definition 1) is no smaller than some $\gamma_{k}>0$, that is
$\gamma_{k}\leq\min_{t\geq t_{0}}\gamma_{k}(x;\theta(t))\wedge\min_{t\geq t_{0}}\gamma_{k}(x^{\prime};\theta^{\prime}(t))$
we have
$\left\langle h_{k}(x;\,\theta(t)),h_{k}(x^{\prime};\,\theta^{\prime}(t))\right\rangle\leq\left\langle h_{k}(x;\,\theta(t_{0})),h_{k}(x^{\prime};\,\theta^{\prime}(t_{0}))\right\rangle\exp\left(-2\lambda(t-t_{0})+2\,\frac{\sqrt{C\,(t-t_{0})}}{\gamma_{k}}\right)$
where $C$ is a constant such that $L(\theta(t_{0}))\vee L^{\prime}(\theta^{\prime}(t_{0}^{\prime}))\leq C$. [Proof]
In particular, this implies
Corollary 2.
Under conditions of Lemma 8
$\lim_{t\to\infty}\inf_{\lambda\in\Lambda(t)}\left\langle h_{k}(x;\,\theta(t)),h_{k}(x^{\prime};\,\theta^{\prime}(t))\right\rangle=0$
where $\Lambda(t)=\left\{\lambda\in[0,1]~{}:~{}\lambda>\frac{\sqrt{C/t}}{\gamma_{k}}\right\}$.
Proof of Theorem 2.
For any given layer $k$, our first step is to show that the pre-activation of a merged model satisfies:
$(W_{k}+W_{k}^{\prime})h_{k-1}(x^{\prime};\,\theta+\theta^{\prime})=W_{k}^{\prime}h_{k-1}(x^{\prime};\,\theta+\theta^{\prime})~{}.$
Assume induction hypothesis on $k$:
$\left\langle h_{k-1}(x_{i}^{\prime};\,\theta^{\prime}),h_{k-1}(x;\,\theta+\theta^{\prime})\right\rangle=0$
that is that activations of different networks on point from respective task, are orthogonal.
Proposition 2 (third statement) gives that for the stationary point
$W_{k}=-\frac{1}{\lambda}\nabla L(W_{k})~{}.$
Then, using this fact we have
$(W_{k}+W_{k}^{\prime})h_{k-1}(x;\,\theta+\theta^{\prime})$
$\quad=-\frac{1}{n\lambda}\sum_{i=1}^{n}y_{i}\,\ell^{\prime}(y_{i}f_{\theta}(x_{i}))\,p_{k,i}\left\langle h_{k-1}(x_{i};\,\theta),h_{k-1}(x;\,\theta+\theta^{\prime})\right\rangle$
$\qquad-\frac{1}{n\lambda}\sum_{i=1}^{n}y_{i}^{\prime}\,\ell^{\prime}(y_{i}^{\prime}f_{\theta^{\prime}}(x_{i}^{\prime}))\,p_{k,i}^{\prime}\left\langle h_{k-1}(x_{i}^{\prime};\,\theta^{\prime}),h_{k-1}(x;\,\theta+\theta^{\prime})\right\rangle$
=$\quad\stackrel{{\scriptstyle(a)}}{{=}}-\frac{1}{n\lambda}\sum_{i=1}^{n}y_{i}\,\ell^{\prime}(y_{i}f_{\theta}(x_{i}))\,p_{k,i}\left\langle h_{k-1}(x_{i};\,\theta),h_{k-1}(x;\,\theta+\theta^{\prime})\right\rangle$
$\quad=W_{k}h_{k-1}(x;\,\theta+\theta^{\prime})$
where in $(a)$ we used in induction hypothesis. Taking $(\cdot)_{+}$ on both sides we have
$h_{k}(x;\,\theta+\theta^{\prime})=(W_{k}h_{k-1}(x;\,\theta+\theta^{\prime}))_{+}$
and expanding this recursion we have
$h_{k}(x;\,\theta+\theta^{\prime})=(W_{k}(W_{k-1}(\dots W_{1}h_{0}(x;\,\theta+\theta^{\prime}))_{+})_{+})_{+}=h_{k}(x;\,\theta)$
since $h_{0}(x;\,\theta+\theta^{\prime})=x$.
Now we show the step of induction, using the above combined with Corollary 2, namely
$\left\langle h_{k}(x_{i}^{\prime};\,\theta^{\prime}),h_{k}(x;\,\theta+\theta^{\prime})\right\rangle=\left\langle h_{k}(x_{i}^{\prime};\,\theta^{\prime}),h_{k}(x;\,\theta)\right\rangle=0~{}.$
Base case of induction is evident since $\left\langle h_{0}(x_{i}^{\prime};\,\theta^{\prime}),h_{0}(x;\,\theta+\theta^{\prime})\right\rangle=\left\langle x_{i}^{\prime},x\right\rangle=0$. ∎
B.5 Proof of Lemma 8
Throughout the proof notation $(h)_{j}$ means that we are considering $j$th coordinate of activation vector $h$, and similarly for $h^{\prime}$. We also use abbreviations
$h(t)$ $=h(x;\,W_{k}(t))$
$h_{k-1}(t)$ $=h(x;\,W_{k-1}(t))$
$h_{i}(t)$ $=h(x_{i};\,W_{k}(t))$
$h_{k-1,i}(t)$ $=h(x_{i};\,W_{k-1}(t))$
$\nabla(h)_{j}(W(t))$ $=\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}W}(h)_{j}(x;\,W)\Bigr{|}_{W=W(t)}$
and similarly for $h^{\prime}$. Observe that
$\nabla(h)_{j}(W(t))=\mathbbm{1}\left((h)_{j}(t)>0\right)\,e_{j}\,h_{k-1}(t)^{\top}~{}.$
In particular the above implies that $\left\langle W(t),\nabla(h)_{j}(W(t))\right\rangle=(h)_{j}(t)$.
Consider the following dynamics of the inner product between activations, where by chain rule we have
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\left\langle h(t),h^{\prime}(t)\right\rangle$ $=\left\langle\dot{h}(t),h^{\prime}(t)\right\rangle+\left\langle h(t),\dot{h}^{\prime}(t)\right\rangle~{}.$
Focusing on one of the terms on the right hand side (another one is handled similarly):
$\left\langle\dot{h}(t),h^{\prime}(t)\right\rangle$ $=\sum_{j}\left\langle\dot{W}(t),\nabla(h)_{j}(W(t))\right\rangle\cdot(h^{\prime})_{j}(t)$
$=-\lambda\sum_{j}\left\langle W(t),\nabla(h)_{j}(W(t))\right\rangle\cdot(h^{\prime})_{j}(t)-\sum_{j}\left\langle\nabla L_{W}(\theta(t)),\nabla(h)_{j}(W(t))\right\rangle\cdot(h^{\prime})_{j}(t)$
$=-\lambda\left\langle h(t),h^{\prime}(t)\right\rangle+a(t)$
where
$a(t)$ $:=-\sum_{j}\left\langle\nabla L_{W}(\theta(t)),\mathbbm{1}\left((h)_{j}(t)>0\right)\cdot(h^{\prime})_{j}(t)\,e_{j}\,h_{k-1}(t)^{\top}\right\rangle$
$\leq\|\nabla L_{W}(\theta(t))\|\,\|h_{k-1}(t)\|\sum_{j=1}^{m}\mathbbm{1}\left((h)_{j}(t)>0\right)\cdot(h^{\prime})_{j}(t)$
$=\|\nabla L_{W}(\theta(t))\|\,\|h_{k-1}(t)\|\sum_{j~{}:~{}(h)_{j}(t)>0}\frac{1}{(h)_{j}(t)}\,(h)_{j}(t)\cdot(h^{\prime})_{j}(t)$
$\leq\frac{1}{\gamma_{k}}\,\|\nabla L_{W}(\theta(t))\|\left\langle h(t),h^{\prime}(t)\right\rangle$
where we recall that $\gamma_{k}\leq\min\left\{\left\langle e_{j},W(t)h_{k-1}(t)\right\rangle/\|h_{k-1}(t)\|~{}:~{}\left\langle e_{j},W(t)h_{k-1}(t)\right\rangle>0~{},t\geq t_{0}~{},j\in[m]\right\}$.
Putting everything together, we have differential inequality
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\left\langle\dot{h}(t),h^{\prime}(t)\right\rangle\leq-2\lambda\left\langle h(t),h^{\prime}(t)\right\rangle+\left(\frac{1}{\gamma_{k}}\,\|\nabla_{W}L(\theta(t))\|+\frac{1}{\gamma^{\prime}_{k}}\,\|\nabla_{W^{\prime}}L^{\prime}(\theta^{\prime}(t))\|\right)\left\langle h(t),h^{\prime}(t)\right\rangle$
and so by Grönwall’s inequality
$\left\langle h(t),h^{\prime}(t)\right\rangle\leq\left\langle h(t_{0}),h^{\prime}(t_{0})\right\rangle\exp\left(-2\lambda(t-t_{0})+\frac{1}{\gamma_{k}}\int_{t_{0}}^{t}\|\nabla_{W}L(\theta(s))\|\mathop{}\!\mathrm{d}s+\frac{1}{\gamma^{\prime}_{k}}\int_{t_{0}}^{t}\|\nabla_{W^{\prime}}L^{\prime}(\theta^{\prime}(s))\|\mathop{}\!\mathrm{d}s\right)~{}.$
Here, the second result of Lemma 7 combined with Jensen’s inequality gives
$\int_{t_{0}}^{t}\|\nabla_{W}L(\theta(s))\|\mathop{}\!\mathrm{d}s\leq\sqrt{(t-t_{0})(L(\theta(t_{0}))-L(\theta(t)))}$
which completes the proof.
Appendix C Additional empirical results from Section 6
C.1 Transformer results
Figure 4 summarizes more measurements in addition to those in Section 6.
Model name number of layers model dimension FF dimension
BERT 12 768 3072
RoBERTa 12 768 3072
GPT2-Large 36 1280 5120
GPT2-XL 48 1600 6400
Phi-2 32 2560 10240
GPT-J 28 4096 16384
Table 1: Specification of Transformer models.
In Table 1 model dimension refers to dimensionality of self-attention weight matrices, while FF dimension refers to the size of the hidden layer of the feedforward network.
Figure 4: Stable ranks and Frobenius norms of different weight matrices in pretrained BERT (first row), GPT2-Large (second row), RoBERTa (third row), and GPT2-XL (fourth row) model.
C.2 MLP results
The second set of experiments is performed on ‘Fashion MNIST’ dataset (Xiao et al., 2017), which contains grayscale images of 28x28 pixels each, representing clothing items from 10 different categories. We adapt this dataset with sample size $10$ for binary classification by grouping first 5 classes into class 0, and remaining into class 1. When we consider two different tasks, we append 784-dimensional zero vector for the first task, and prepend in case of the second task. This way, inputs from different tasks remain orthogonal, while the length of each input is preserved. Finally, in case of task one binary labels are preserved as is, while for the second task labels are inverted.
Figure 5: (Fashion MNIST dataset) First row: Training neural networks on different tasks (orthogonal inputs) (left) vs. the same task (right) and merging the parameters by adding weight matrices. The resulting network performs well on different tasks after sufficiently many iterations, while given the same task, it does not. Second row: this effect manifests when weight decay strength is sufficiently large (left). Stable rank of each weight matrix converges to a small value. Merged network matches the stable rank of individual networks (right).